16 if x/2+y/5=1,solve for y. Please show all steps.
Answer by checkley75(3666)

x/2+y/5=1
(5x+2y)/10=1
5x+2y=10
2y=-5x+10
y=-5x/2+10/2
y=-2.5x+5